OpenStudy (anonymous):

definite integral (tan^-1 x) / (x^2 + 1). upper bound = infinity, lower bound = 1. use improper integral formula. show steps.

OpenStudy (anonymous):

\[ u =\tan^{-1} x \\ du = \frac {dx}{1+x^2}\\ \int_{\frac \pi 4}^{\frac \pi 2} u du=\frac{3 \pi ^2}{32} \]

OpenStudy (anonymous):

how'd you get those bound values?

OpenStudy (eyust707):

Initially the integral was: \[\int\limits_{x=1}^{x=\infty} ... \] But then you changed to from x's to u's so we also had to change the bounds Since \(u = \tan^{-1} x \) we just plug in each bound to that equation to get the new bounds of the integral. \[\tan^{-1} \infty = {\pi \over 2}\] \[\tan^{-1} 1 = {\pi \over 4}\]

OpenStudy (anonymous):

hmm, yes, that much makes sense.

OpenStudy (anonymous):

Do you need any more details?

OpenStudy (anonymous):

no, that's good thanks